# Mechanisms for producing Primordial Black Holes from Inflationary Models Beyond Fine-Tuning

### v1

@@ Abstract
In this study, we present an analysis of the fine-tuning required in various inflationary models in order to explain the production of Primordial Black Holes (PBHs). We specifically examine the degree of fine-tuning necessary in two prominent single field inflationary models: those with an inflection point and those with step-like features in the potential. Our findings indicate that models with step-like features generally require less fine-tuning compared to those with an inflection point, making them more viable for consistent PBH production. An interesting outcome of these models is that, in addition to improved fine-tuning, they may also predict low-frequency signals that can be detected by pulsar timing array (PTA) collaborations. Additionally, we extend our analysis to multifield inflationary models to assess whether the integration of additional fields can further alleviate the fine-tuning demands. The study also explores the role of a spectator field and its impact on the fine-tuning process. Through a comparative analysis across these models, we evaluate their parametric sensitivities and alignment with the constraints from Cosmic Microwave Background (CMB) observations. Our results indicate that although mechanisms involving a spectator field can circumvent the issue of fine-tuning parameters for PBH production, both multifield models and models with step-like features present promising alternatives. While fine-tuning involves multiple considerations, our primary objective is to evaluate various inflationary models to identify the one that most naturally explains the formation of PBHs. Hence, this study aims to the development of less constrained inflationary scenarios that align with observational data and support the theoretical possibility of PBH production.
@@ 1 Introduction
In recent years, there has been an increasing focus on primordial black holes (PBHs) due to their possible contribution to explain the dark matter of the Universe. This explanation has gained significant attention following the detection of gravitational waves (GWs) emitted by binary black hole mergers, as reported by the LIGO/VIRGO/Kagra (LVK) collaboration [1, 2, 3, 4, 5]. Moreover, the recent detection of GWs at nano-Hertz frequencies by pulsar timing array (PTA) collaborations [6, 7, 8, 9, 10, 11, 12] has sparked additional interest in the potential link between these low-frequency GWs and PBHs . Consequently, numerous theoretical studies have explored the potential link between dark matter, PBHs and GWs [13, 14, 15, 16, 17, 18, 19]. All these studies must adhere to the observational constraints on inflation from Cosmic Microwave Background (CMB) anisotropies as reported by the Planck collaboration [20, 21].
@@
In previous theoretical researches, it is suggested that PBHs can form through a significant enhancement of the curvature perturbations at small scales. One widely adopted mechanism for producing PBHs from inflation involves studying an inflection point in the effective scalar potential [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Such inflationary potentials are generally characterized by a slow-roll phase at large scales exit the horizon, followed by an inflection point region at small scales where the power spectrum undergoes a significant enhancement. However, a major drawback of this mechanism is the high level of fine-tuning required for the parameters to achieve such a configuration in the potential [27, 23, 37].
@@
Alternative mechanisms within the framework of single field inflation have also been proposed to alleviate the problematic fine-tuning [38, 39, 40, 41, 42, 43, 44, 45, 46, 47]. Mechanisms involving step-like features in potentials have been studied [38, 40, 41, 42]. These mechanisms can lead to a significant enhancement of the power spectrum, which is capable of explaining the production of PBHs. Interestingly, this mechanism predicts additional oscillatory patterns in the power spectra, that may lead to a characteristic signal in the induced GWs [48].
@@
The study of an amplification of scalar power spectrum has been extend also to multi-field models. Multi-field inflationary models have been widely explored for the production of PBHs [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. Models with a non-canonical kinetic term in the Lagrangian can lead to a significant enhancement in the scalar power spectrum [53, 54, 55, 56, 57, 58, 59]. Moreover, multi-field models within the framework of natural inflation, such as those involving multiple axion fields [50, 51], have been shown to result in a significant fractional abundance of PBHs. Finally, hybrid models with a waterfall trajectory are acquired a lot of interest as they provide a more natural explanation to PBHs [60, 61, 62, 63, 64, 65].
@@
Furthermore, mechanisms involving a spectator field have been extensively explored as promising alternatives to mitigate the fine-tuning required for PBH production [66, 67, 68, 69, 70, 71, 72, 73]. Although the spectator field does not drive inflation, it evolves independently and significantly influences the formation of structures and phenomena, including PBHs. Notably, it has been demonstrated that these models can circumvent the need for fine-tuning the potential’s parameters to facilitate PBH production [67, 68].
@@
This paper explores a diverse array of mechanisms proposed for the production of PBHs, as studied in recent literature. First, we delve into single field inflation models, examining both the inflection point mechanism and models that feature step-like characteristics, while evaluating the associated fine-tuning demands. We found that the amount of fine-tuning required for models with an inflection point is significantly large, whereas the step-like mechanism can lead to a more natural explanation as the amount of fine-tuning is significantly reduced. Notably, this mechanism can explain GW signals at low frequencies. We discuss this implementation in a separated section and leave the details for a future paper.
@@
In the continuation of our analysis, we extend our study to multifield inflationary models. Specifically, we explore two-field model mechanisms: the first incorporates non-canonical kinetic terms in the Lagrangian, and the second features a waterfall trajectory typical of hybrid inflationary models. We find that the amount of fine-tuning is significantly reduced in both multifield mechanisms by at least three orders of magnitude in comparison with inflection point models. Lastly, we discuss the implications of models that include a spectator field for explaining the production of PBHs, highlighting that the need for fine-tuning to describe such structures can be avoided.
@@
The layout of this paper is as follows: In Section 2, we describe our methodology for computing the amount of fine-tuning. In Section 3, we present an analysis of fine-tuning in single field inflationary mechanisms. In Section 4, we extend this analysis to two-field models. In Section 5, we discuss the amount of fine-tuning in models with a spectator field. In Section 6, we explore an interesting application of step-like potentials to scalar-induced GWs. Finally, in Section 7, we present our conclusions.
@@ 2 Fine-tuning
To assess the fine-tuning of parameters, we employ the quantity $\Delta_{p_{i}}$, defined as the maximum value of the following logarithmic derivative [74]:
@@
3 Single field models
In this section, we discuss single field inflationary models for the production of PBHs. These models are widely recognized in the literature for their role in the formation of PBHs and offer several advantages. First, they provide a simple mechanism for producing PBHs during inflation. Second, they can explain a significant fraction of dark matter. Third, they are easily adaptable to various theoretical frameworks.
In the subsequent sections, we will explore two distinct mechanisms. The first involves a mechanism characterized by an inflection point in the scalar potential [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. The second examines models that incorporate step-like features [40, 41]. We compute the amount of fine-tuning for both these two different mechanism. Our result suggest that models with step-like can significantly reduce this required tuning of the parameters.
@@ 3.1.1 Description of the mechanism
where $\phi_{\text{inf}}$ denotes the inflection point. At this point, the velocity of the inflaton field significantly decreases, leading to a local enhancement while the inflaton field remains almost constant. During this plateau, the power spectrum experiences significant enhancement, facilitating the production of PBHs. In this work, we also include in this category models with bump features in the potential, such as those in [75, 33, 76], since they satisfy the conditions given in Eq.(2).
As we mentioned before, the concept of an inflection point in the potential has been widely adopted in theoretical studies. The degree of fine-tuning required for inflection point models has been extensively analyzed, notably in Ref. [37]. According to this study, polynomial potentials appear to require substantial fine-tuning both to explain the formation of PBHs and to meet the observational constraints of inflation at CMB scales.
@@ 3.1.2 The issue of fine-tuning
Among the various models proposing an inflection point, we focus on the one detailed in Ref. [22]. This model is also highlighted in Ref. [37] as one of the models requiring minimal fine-tuning within the inflection point framework. The potential, as described in Ref. [22], is given by:
@@
$a=1/\sqrt{2},\quad b=2/3,\quad\lambda=1.86\times 10^{-6},\quad\upsilon=0.19669$ (4)
and for initial condition we assume $\phi_{ic}=2.4719$  [^1] We assume always that the initial conditions are taken at the pivot scales ($\rm{k=0.05M_{pc}^{-1}}$) in order to have a prompt comparison with the observable constraints of inflation. This choice of initial condition lead to the following prediction of spectral index, $\rm{n_{s}}$, and tensor-to-scalar ration, $\rm{r}$:
@@
In Fig. 1, we depict the power spectrum for a small variation of the parameter $b$, which influences the inflection point and consequently requires more fine-tuning. This figure clearly demonstrates that the peak value is highly sensitive to variations in this parameter.
Figure 1: The power spectrum derived from the potential described in Eq. (3), showing the effects of small variations in the parameter b of order $\mathcal{O}(10^{-6})$.
To systematically quantify the fine-tuning, we evaluate the quantity of Eq. (1).  [^2] The evalution of power spectrum is briefly described in Appendix A.1. In Table 1, we present the evaluation of Eq. (1) for the potential (3). Our results are consistent with those proposed by Cole et al. [37] in their study of fine-tuning for the peak of the scalar power spectrum. The parameter $\rm b$ can be given by the parameter $\rm a$ by the following expression[22, 37]:
@@
This parameter defines the inflection point. In contrast to [37], we also evaluate the fine-tuning of the parameter $\rm b$ individually, raising questions about the necessity of redefining this function and uncovering hidden fine-tuning and we found that this parameter need a lot of fine-tuning. As discussed in [37], the amount of fine-tuning required for models with an inflection point is significant; with variations in the studied models, it can reach an order of magnitude as high as $\rm{10^{8}}$. The fine-tuning associated with models featuring an inflection point has also been studied in [27, 28].
@@
Table 1: Analysis of fine-tuning for the potential given in Eq. (3).
In our analysis, we also evaluate the degree of fine-tuning required for the fractional abundance of PBHs, denoted as $f\rm{{}_{PBH}}$. So, we compute the $\Delta_{p_{i}}^{f_{\rm{PBH}}}$ from Eq.(1). We found that 333 The evaluation of fractional abundance of PBHs, $f_{\rm{PBH}}$, is given in Appendix A.2.
@@
which indicates that achieving the desired fractional abundance necessitates fine-tuning that is one-two order of magnitudes more stringent than that required for the parameters outlined in Table 1. This elevated level of fine-tuning underscores the sensitivity of PBH formation to the specific dynamics of the inflationary model parameters.

### v2

# primordial black holes; inflation; dark matter; scalar induced gravitational waves

@@ Abstract
In this study, we present an analysis of the fine-tuning required in various inflationary models in order to explain the production of Primordial Black Holes (PBHs). We specifically examine the degree of fine-tuning necessary in two prominent single-field inflationary models: those with an inflection point and those with step-like features in the potential. Our findings indicate that models with step-like features generally require less fine-tuning compared to those with an inflection point, making them more viable for consistent PBH production. An interesting outcome of these models is that, in addition to improved fine-tuning, they may also predict low-frequency signals that can be detected by pulsar timing array (PTA) collaborations. Additionally, we extend our analysis to multifield inflationary models to assess whether the integration of additional fields can further alleviate the fine-tuning demands. The study also explores the role of a spectator field and its impact on the fine-tuning process. Our results indicate that although mechanisms involving a spectator field can circumvent the issue of fine-tuning parameters for PBH production, both multifield models and models with step-like features present promising alternatives. While fine-tuning involves multiple considerations, our primary objective is to evaluate various inflationary models to identify the one that most naturally explains the formation of PBHs. Hence, this study introduces a novel approach by categorizing existing PBH mechanisms, paving the way for subsequent research to prioritize models that minimize the need for extensive fine-tuning.
keywords:
primordial black holes; inflation; dark matter; scalar induced gravitational waves
[…]
@@ 1 Introduction
In recent years, there has been an increasing focus on primordial black holes (PBHs) due to their possible contribution to explaining the dark matter of the Universe. This explanation has gained significant attention following the detection of gravitational waves (GWs) emitted by binary black hole mergers, as reported by the LIGO/VIRGO/Kagra (LVK) collaboration Abbott et al. (2016a, b, 2017a, 2017b, 2017c). Moreover, the recent detection of GWs at nano-Hertz frequencies by pulsar timing array (PTA) collaborations Agazie et al. (2023a, b); Afzal et al. (2023); Antoniadis et al. (2023a, b, c, d) has sparked additional interest in the potential link between these low-frequency GWs and PBHs . Consequently, numerous theoretical studies have explored the potential link between dark matter, PBHs, and GWs Özsoy and Tasinato (2023); Sasaki et al. (2018); Carr et al. (2024); Khlopov (2010); Kuhnel and Stamou (2024); Ellis et al. (2024); Vaskonen and Veermäe (2021); Escrivà et al. (2022). All these studies must adhere to the observational constraints on inflation from Cosmic Microwave Background (CMB) anisotropies as reported by the Planck collaboration Ade et al. (2016); Akrami et al. (2020).
@@
In previous theoretical research, it has been suggested that PBHs can form through a significant enhancement of curvature perturbations at small scales. One widely adopted mechanism for producing PBHs from inflation involves studying an inflection point in the effective scalar potential Garcia-Bellido and Ruiz Morales (2017); Hertzberg and Yamada (2018); Ballesteros and Taoso (2018); Ballesteros et al. (2020); Cicoli et al. (2018); Stamou (2021); Spanos and Stamou (2023); Dalianis et al. (2019); Germani and Prokopec (2017); Ezquiaga et al. (2018); Özsoy et al. (2018); Özsoy and Lalak (2021); Karam et al. (2023); Aldabergenov and Ketov (2023); Kannike et al. (2017); Kawasaki et al. (2016); Choudhury and Mazumdar (2014). Such inflationary potentials are generally characterized by a slow-roll phase at large scales exiting the horizon, followed by an inflection point region at small scales where the power spectrum undergoes a significant enhancement. However, a major drawback of this mechanism is the high level of fine-tuning that the parameters require to achieve such a configuration in the potential Stamou (2021); Hertzberg and Yamada (2018); Cole et al. (2023). It is important to note that, within this work, ‘fine-tuning’ refers to the necessity for highly precise parameters to facilitate the abundant production of PBHs. A model is deemed fine-tuned if minor variations in parameters significantly influence the enhancement of the power spectrum and, consequently, the production of PBHs.
@@
Alternative mechanisms within the framework of single-field inflation have also been proposed to alleviate problematic fine-tuning Inomata et al. (2022); Cotner and Kusenko (2017); Kefala et al. (2021); Dalianis et al. (2021); Cai et al. (2022); Choudhury et al. (2023, 2024); Gu et al. (2023); Drees and Erfani (2012); Domènech et al. (2024); Caravano et al. (2024); Sharma et al. (2024); Dimastrogiovanni et al. (2024); Karam et al. (2023). Mechanisms involving step-like features in potentials have been studied Inomata et al. (2022); Kefala et al. (2021); Dalianis et al. (2021); Cai et al. (2022). These mechanisms can lead to a significant enhancement of the power spectrum, which is capable of explaining the production of PBHs. Interestingly, this mechanism predicts additional oscillatory patterns in the power spectra that may lead to a characteristic signal in the induced GWs Fumagalli et al. (2021).
@@
The study of an amplification of scalar power spectrum has also been extended to multi-field models. Multi-field inflationary models have been widely explored for the production of PBHs Palma et al. (2020); Anguelova (2021); Iacconi and Mulryne (2023); Zhou et al. (2020); Mavromatos et al. (2022); Kawai and Kim (2023); Aldabergenov et al. (2022); Braglia et al. (2020); Pi and Sasaki (2023); Gundhi et al. (2021); Geller et al. (2022); Chen et al. (2023); Wang et al. (2024). Models with a non-canonical kinetic term in the Lagrangian can lead to a significant enhancement in the scalar power spectrum Aldabergenov et al. (2022); Braglia et al. (2020); Pi and Sasaki (2023); Gundhi et al. (2021); Geller et al. (2022); Chen et al. (2023); Wang et al. (2024); Heydari and Karami (2022); Pi et al. (2018). Moreover, multi-field models within the framework of natural inflation, such as those involving multiple axion fields Zhou et al. (2020); Mavromatos et al. (2022), have been shown to result in a significant fractional abundance of PBHs. Finally, hybrid models with a waterfall trajectory have acquired a lot of interest as they provide a more natural explanation to PBHs Afzal and Ghoshal (2024); Spanos and Stamou (2021); Braglia et al. (2023); Clesse and García-Bellido (2015); Tada and Yamada (2023); Dimopoulos (2022).
@@
Furthermore, mechanisms involving a spectator field have been extensively explored as promising alternatives to mitigate the fine-tuning required for PBH production Carr et al. (2017); Kohri et al. (2013); Stamou and Clesse (2024, 2023); Stamou (2024); Cai et al. (2021); Gow et al. (2023); Hooper et al. (2024). Although the spectator field does not drive inflation, it evolves independently and significantly influences the formation of structures and phenomena, including PBHs. Notably, it has been demonstrated that these models can circumvent the need for fine-tuning the potential’s parameters to facilitate PBH production Kohri et al. (2013); Stamou and Clesse (2024).
This paper investigates the extent of fine-tuning required across a diverse range of mechanisms previously proposed in the literature. Specifically, we examine four characteristic inflationary mechanisms for producing primordial black holes (PBHs): a single-field mechanism with an inflection point, one with a step-like feature, a two-field mechanism with a non-canonical kinetic term, and a mechanism incorporating a waterfall trajectory. For each PBH production mechanism, we provide a rationale for selecting a particular model. This study diverges from the findings reported in Ref. Cole et al. (2023) as it does not solely concentrate on single-field inflationary models with an inflection point or bulk in their scalar potential. While acknowledging the contributions of these mechanisms, our analysis broadens to include additional inflationary models, such as those involving multiple fields, as discussed earlier.
@@
First, we delve into single-field inflation models, examining both the inflection point mechanism and models that feature step-like characteristics while evaluating the associated fine-tuning demands. We found that the amount of fine-tuning required for models with an inflection point is significantly large, whereas the step-like mechanism can lead to a more natural explanation as the amount of fine-tuning is significantly reduced. Notably, this mechanism can explain GW signals at low frequencies. We discuss this implementation in a separated section and leave the details for a future paper.
@@
In the continuation of our analysis, we extend our study to analyze the fine-tuning of multifield inflationary models. Specifically, we explore two-field model mechanisms: the first incorporates non-canonical kinetic terms in the Lagrangian, and the second features a waterfall trajectory typical of hybrid inflationary models. We find that the amount of fine-tuning is significantly reduced in both multifield mechanisms by at least three orders of magnitude in comparison with inflection point models. Lastly, we discuss the implications of models that include a spectator field for explaining the production of PBHs, highlighting that the need for fine-tuning to describe such structures can be avoided. Therefore, the novelty of this work lies in classifying the proposed PBH mechanisms, focusing on determining which mechanisms can more naturally explain the production of PBHs. This approach can contribute to future investigations towards the mechanisms that inherently require less fine-tuning.
@@
The layout of this paper is as follows: In Section 2, we describe our methodology for computing the amount of fine-tuning. In Section 3, we present an analysis of fine-tuning in single-field inflationary mechanisms. In Section 4, we extend this analysis to two-field models. In Section 5, we discuss the amount of fine-tuning in models with a spectator field. In Section 6, we explore an interesting application of step-like potentials to scalar-induced GWs. Finally, in Section 7, we present our conclusions. The Appendices provide the detailed equations that are needed for our analysis: Appendix A covers the evaluation of the scalar power spectrum, Appendix B details the fractional abundances of PBHs, and Appendix C discusses the energy densities of GWs.
@@ 2 Fine-Tuning
To assess the fine-tuning of parameters, we employ the quantity $\Delta_{p_{i}}$, defined as the maximum value of the following logarithmic derivative Barbieri and Giudice (1988):
@@
3 Single-Field Models
In this section, we discuss single-field inflationary models for the production of PBHs. These models are widely recognized in the literature for their role in the formation of PBHs and offer several advantages. First, they provide a simple mechanism for producing PBHs during inflation. Second, they can explain a significant fraction of dark matter. Third, they are easily adaptable to various theoretical frameworks.
In the subsequent sections, we will explore two distinct mechanisms. The first involves a mechanism characterized by an inflection point in the scalar potential Garcia-Bellido and Ruiz Morales (2017); Hertzberg and Yamada (2018); Ballesteros and Taoso (2018); Ballesteros et al. (2020); Cicoli et al. (2018); Stamou (2021); Spanos and Stamou (2023); Dalianis et al. (2019); Germani and Prokopec (2017); Ezquiaga et al. (2018); Özsoy et al. (2018); Özsoy and Lalak (2021); Karam et al. (2023); Aldabergenov and Ketov (2023); Kannike et al. (2017). The second examines models that incorporate step-like features Kefala et al. (2021); Dalianis et al. (2021). We compute the amount of fine-tuning for both of these two different mechanisms. Our results suggest that models with step-like features can significantly reduce this required tuning of the parameters.
@@ 3.1.1 Description of the Mechanism
where $\phi_{\text{inf}}$ denotes the inflection point. At this point, the velocity of the inflaton field significantly decreases, leading to a local enhancement while the inflaton field remains almost constant. During this plateau, the power spectrum experiences significant enhancement, facilitating the production of PBHs. In this work, we also include in this category models with bump features in the potential, such as those in Mishra and Sahni (2020); Özsoy and Lalak (2021); Fu and Chen (2023), since they satisfy the conditions given in Equation (2).
As we mentioned before, the concept of an inflection point in the potential has been widely adopted in theoretical studies. The degree of fine-tuning required for inflection point models has been extensively analyzed, notably in Ref. Cole et al. (2023). According to this study, polynomial potentials appear to require substantial fine-tuning both to explain the formation of PBHs and to meet the observational constraints of inflation at CMB scales.
@@
Among the various models proposing an inflection point, we focus on the one detailed in Ref. Garcia-Bellido and Ruiz Morales (2017). This model is also highlighted in Ref. Cole et al. (2023) as one of the models requiring minimal fine-tuning within the inflection point framework. The potential, as described in Ref. Garcia-Bellido and Ruiz Morales (2017), is given by:
@@
$a=1/\sqrt{2},\quad b={3/2},\quad\lambda=1.86\times 10^{-6},\quad\upsilon=0.19669$ (4)
and for initial condition we assume $\phi_{ic}=2.4719$ (we assume always that the initial conditions are taken at the pivot scales ($\rm{k=0.05M_{pc}^{-1}}$) in order to make a prompt comparison with the observable constraints of inflation). This choice of initial condition led to the following prediction of the spectral index, $\rm{n_{s}}$, and the tensor-to-scalar ration, $\rm{r}$ (for the evaluation of spectral index $n_{s}$ and $r$, we assume that $n_{s}\simeq 1+2\eta_{V}-6\varepsilon_{V}$ and $r\simeq 16\varepsilon_{V}$, where $\varepsilon_{V}=\frac{1}{2}\left({V^{\prime}(\phi)}/{V(\phi)}\right)^{2},\quad\eta_{V}={V^{\prime\prime}(\phi)}/{V(\phi)}$.):
@@
In Figure 1, we depict the power spectrum for a small variation of the parameter $b$, which influences the inflection point and consequently requires more fine-tuning. This figure clearly demonstrates that the peak value is highly sensitive to variations in this parameter.
Figure 1: The power spectrum derived from the potential described in Equation (3), showing the effects of small variations in the parameter b of order $\mathcal{O}(10^{-6})$. The purple line corresponds to the parameters given in Equation (4).
To systematically quantify the fine-tuning, we evaluate the quantity of Equation (1). (The evalution of power spectrum is briefly described in Appendix A.) In Table 1, we present the evaluation of Equation (1) for the potential (3). Our results are consistent with those proposed by Cole et al. Cole et al. (2023) in their study of fine-tuning for the peak of the scalar power spectrum. The parameter $\rm b$ can be given by the parameter $\rm a$ by the following expression Garcia-Bellido and Ruiz Morales (2017); Cole et al. (2023):
@@
This parameter defines the inflection point. In contrast to Cole et al. (2023), we also evaluate the fine-tuning of the parameter $b$ individually, raising questions about the necessity of redefining this function and uncovering hidden fine-tuning, and we found that this parameter need a lot of fine-tuning. As discussed in Cole et al. (2023), the amount of fine-tuning required for models with an inflection point is significant; with variations in the studied models, it can reach an order of magnitude as high as $\rm{10^{8}}$. The fine-tuning associated with models featuring an inflection point has also been studied in Stamou (2021); Spanos and Stamou (2023).
Table 1: Analysis of fine-tuning for the potential given in Equation (3).
@@
In our analysis, we also evaluate the degree of fine-tuning required for the fractional abundance of PBHs, denoted as $f\rm{{}_{PBH}}$. So, we compute the $\Delta_{p_{i}}^{f_{\rm{PBH}}}$ from Equation (1). We found that
@@
which indicates that achieving the desired fractional abundance necessitates fine-tuning that is 1–2 orders of magnitude more stringent than that required for the parameters outlined in Table 1. This elevated level of fine-tuning underscores the sensitivity of PBH formation to the specific dynamics of the inflationary model parameters. The evaluation of the fractional abundance of PBHs, $f_{\text{PBH}}$, is detailed in Appendix B. In this Appendix, we also present the fractional abundances of PBHs for all the mechanisms discussed in this work.